### v1

@@ Abstract
We introduce a privacy auditing scheme for ML models that relies on membership inference attacks using generated data as “non-members”. This scheme, which we call PANORAMIA, quantifies the privacy leakage for large-scale ML models without control of the training process or model re-training and only requires access to a subset of the training data. To demonstrate its applicability, we evaluate our auditing scheme across multiple ML domains, ranging from image and tabular data classification to large-scale language models.
@@ 1 Introduction
Machine learning (ML) models have become increasingly prevalent in our society, including in applications such as healthcare, finance, or social networks. As a result, ML models are often trained on personal and private data, and maintaining the privacy of this data is a major concern. Unfortunately, access to ML models, and even their predictions, leaks information about their training data. In particular, ML models are susceptible including Membership Inference Attacks (MIA) (e.g., inferring that a particular profile was part of the training set for a disease detection model) Shokri et al. (2017); Choquette-Choo et al. (2020); Song & Mittal (2020); Carlini et al. (2022a) or reconstruction attacks (e.g., recovering social security numbers from a language model) Carlini et al. (2019; 2021); Balle et al. (2022).
Differential Privacy (DP) Dwork et al. (2006) is the leading approach for preventing such data leakage. Training ML models with DP, for instance by using DP-SGD Abadi et al. (2016), upper-bounds the worst-case privacy loss incurred by the training data. Thus, the amount of information that an adversary can gain by observing the model or its predictions is limited. Nonetheless, despite recent advances, DP still decreases the model accuracy even for values of the privacy parameters offering weak theoretical guarantees.
In an effort to reconcile the need for privacy and accuracy in ML deployments, auditing schemes empirically assess the privacy leakage of a target ML model or algorithm Nasr et al. (2021); Jagielski et al. (2020); Zanella-Béguelin et al. (2022); Lu et al. (2023); Nasr et al. (2023); Steinke et al. (2023) whether trained with DP or not. More precisely, auditing schemes use the DP definition and parameters to formalize and quantify privacy leakage. In practice, privacy auditing usually relies on the link between DP and MIA performance Wasserman & Zhou (2010); Kairouz et al. (2015); Dong et al. (2019) to lower-bound the privacy loss of an ML model or algorithm. At a high level, DP implies an upper-bound on the performance of MIAs, and thus empirically violating this bound hence refutes a DP claim. When performing an audit, the stronger the MIA, the tighter the audit, with improvements in MIA performance leading directly to a tighter bound. Auditing schemes have proven valuable in many settings, such as to audit DP implementations Nasr et al. (2023), or to study the tightness of DP algorithms Nasr et al. (2021); Lu et al. (2023); Steinke et al. (2023).
Privacy audits leverage three key primitives for strong MIAs. First, auditors with access to the target model’s data and training pipeline use a large number of shadow models, which are trained models used as examples of behavior in the presence or absence of specific data points, to perform efficient MIAs Carlini et al. (2022a); Zanella-Béguelin et al. (2022); Nasr et al. (2023). Second, auditors with control over the training pipeline rely on canaries, datapoints specially crafted to be easy to detect when added to the training set Nasr et al. (2023); Lu et al. (2023); Steinke et al. (2023). Third, auditors can plug inside the learning algorithm to leverage fine-grained information, by measuring or changing the gradients in DP-SGD Nasr et al. (2021). However, these primitives suffer from severe shortcomings. For instance, approaches based on shadow models are costly, as they require numerous model retraining. Adding canaries requires control of the training data, and leads to polluting the final model, which is better suited to auditing algorithms rather than trained models.
These shortcomings prevent the application of privacy audits in at least three important settings. First, it prevents audits of complex ML pipelines in which the model is split among different actors, such as internal teams owning embeddings and predictive models, or advertising pipelines split between advertisers and publishers. Indeed, in these settings, a single actor cannot control the whole training set or retrain the whole pipeline several times. Second, participants co-training a model with Federated Learning (FL) cannot easily audit the privacy leak of the trained model, released to all participants, with respect to their own data as this would require either polluting the final model with canaries or collaborating honestly with all clients to train shadow models. Third, users of ML-driven applications cannot measure privacy leakage a posteriori, as they cannot retroactively add data and cannot retrain the model. This prevents asking questions such as how much a model trained on one’s data reveals this information.
We propose PANORAMIA, a new scheme for Privacy Auditing with NO Retraining by using Artificial data for Membership Inference Attacks. We build on the key insight that the bottleneck limiting privacy audits’ performance and applicability is the access to examples of both training set data (i.e., members) and comparable data that is not in the training set (i.e., non-members). Indeed, shadow models which are often used to train a strong MIA, are models retrained on different subsets of the training data. On the other hand, canaries introduce new member data points, typically from a different distribution, that are easier to distinguish from non-members and reduce the requirement for a large number of members and non-members.
@@
In PANORAMIA, we consider an auditor with access to a subset of the training data (e.g., the training set of a participant in FL or an ML model user’s own data), and introduce a new alternative for accessing non-members: using synthetic datapoints from a generative model trained on the member data. PANORAMIA uses this generated data, together with known members, to train and evaluate a MIA attack on the target model to audit (§3). We also adapt the theory of privacy audits, and show how PANORAMIA can estimate the privacy loss (though not a lower-bound) of the target model with regards to the known member subset (§4). An important benefit of PANORAMIA is to perform privacy audits with (1) no retraining the target ML model (i.e., we audit the end-model, not the training algorithm), (2) no alteration of the model, dataset, or training procedure and (3) only partial knowledge of the training set. These characteristics make PANORAMIA applicable to all use-cases listed above. To demonstrate this, we evaluate PANORAMIA on three data modalities (images, text, and tabular data), and show that it can estimate significant privacy loss up to $\tilde{\epsilon}=1.28$. Comparing privacy leakage across models, we observe that overfitted models, larger models as well as models with larger DP parameters have higher privacy leakage.
Figure 1: PANORAMIA’s two-phase privacy audit. Phase 1: training of generative model ${\mathcal{G}}$ using member data. Phase 2: training a MIA on a subset of member data and generated non-member data, along with the loss values of $f$ on these data points. The, the comparison of the performance of the MIA to a baseline classifier that does not have access to $f$ is performed. For notation please refer to Table 1.
2 Background
@@
DP is the established privacy definition in the context of ML models as well as data analysis in general. We focus on the pure DP definition to quantify privacy with well-understood semantics. In a nutshell, DP is a property of a randomized mechanism (or computation) from datasets to an output space ${\mathcal{O}}$, noted $M:{\mathcal{D}}\rightarrow{\mathcal{O}}$. It is defined over neighboring datasets $D,D^{\prime}$, that differ by one element $x\in{\mathcal{X}}$ (we use the add/remove neighboring definition), which is $D^{\prime}=D\cup\{x\}$. Formally:
@@
A randomized mechanism $M:{\mathcal{D}}\rightarrow{\mathcal{O}}$ is $\epsilon$-DP if for any two neighboring datasets $D$, $D^{\prime}\in{\mathcal{D}}$, and for any measurable output subset $O\subseteq{\mathcal{O}}$ it holds that:
$P[M(D)\subseteq O]\leq e^{\varepsilon}P[M(D^{\prime})\subseteq O].$
@@
Since the neighboring definition is symmetric, so is the DP definition, and we also have that $P[M(D^{\prime})\subseteq O]\leq e^{\varepsilon}P[M(D)\subseteq O]$. Intuitively, $\epsilon$ upper-bounds the worst-case contribution of any individual example to the distribution over outputs of the computation (i.e., the model learned). More formally, $\epsilon$ is an upper-bound on the privacy loss incurred by observing an output $o$, defined as $\big{|}\ln\left(\frac{{\mathbb{P}}[M(D)=o]}{{\mathbb{P}}[M(D^{\prime})=o]}\right)\big{|}$, which quantifies how much an adversary can learn to distinguish $D$ and $D^{\prime}$ based on observing output $o$ from $M$. A smaller $\epsilon$ hence means higher privacy.
@@
DP, MIA and privacy audits. To audit a DP training algorithm $M$ that outputs a model $f$, one can perform a MIA on datapoint $x$, trying to distinguish between a neighboring training sets $D$ and $D^{\prime}=D\cup\{x\}$. The MIA can be formalized as a hypothesis test to distinguish between ${\mathcal{H}}_{0}=D$ and ${\mathcal{H}}_{1}=D^{\prime}$ using the output of the computation $f$. Wasserman & Zhou (2010); Kairouz et al. (2015); Dong et al. (2019) show that any such test at significance level $\alpha$ (the False Positive Rate (FPR), or probability of rejecting ${\mathcal{H}}_{0}$ when it is true) has power (True Positive Rate (TPR), or the probability of rejecting ${\mathcal{H}}_{0}$ when ${\mathcal{H}}_{1}$ is indeed true) bounded by $e^{\epsilon}\alpha$. In practice, one repeats the process of training model $f$ with and without $x$ in the training set, and uses a MIA to guess whether $x$ was included. If the MIA has $\textrm{TPR}>e^{\epsilon}\textrm{FPR}$, the training procedure that outputs $f$ is not $\epsilon$-DP. This is the building block of most privacy audits Jagielski et al. (2020); Nasr et al. (2021); Zanella-Béguelin et al. (2022); Lu et al. (2023); Nasr et al. (2023).
@@
Averaging over data instead of models with O(1). The above result bounds the success rate of MIAs when performed over several retrained models, on two alternative datasets $D$ and $D^{\prime}$. Steinke et al. (2023) show that it is possible to average over data when several data points independently differ between $D$ and $D^{\prime}$. Let $x_{1,m}$ be the number of data points independently included in the training set, and $s_{1,m}\in\{0,1\}^{m}$ be the vector encoding inclusion. $T_{0,m}\in{\mathbb{R}}^{m}$ represents any vector of guesses, with positive values for inclusion in the training set (member), negative values for non-member, and zero for abstaining. Then, if the training procedure is $\epsilon$-DP, Proposition 5.1 in Steinke et al. (2023) bounds the performance of guesses from $T$ with:
${\mathbb{P}}\Big{[}\sum_{i=1}^{m}$ $\max\{0,T_{i}\cdot S_{i}\}\geq v\ |\ T=t\Big{]}$
@@
$\leq\underset{{S^{\prime}\sim\textrm{Bernoulli}(\frac{e^{\epsilon}}{1+e^{\epsilon}})^{m}}}{{\mathbb{P}}}\Big{[}\sum_{i=1}^{m}|t_{i}|\cdot S^{\prime}_{i}\geq v\Big{]}.$
@@
In other words, an audit (MIA) $T$ that can guess membership better than a Bernoulli random variable with probability $\frac{e^{\epsilon}}{1+e^{\epsilon}}$ refutes $\epsilon$-DP claim. In this work we build on this result, extending the algorithm (§3) and theoretical analysis (§4) to enable the use of generated data for non-members, which breaks the independence between data points.
@@ 3 PANORAMIA
Notation Description
$f$ the target model to be audited.
${\mathcal{G}}$ the generative model
${\mathcal{D}}$ distribution over auditor samples
$D_{f}$ (subset of the) training set of the target model $f$ from ${\mathcal{D}}$
$m$ $|D_{\textrm{in}}|=|D_{\textrm{out}}|\triangleq m$
$b$ baseline classifier for $D_{\textrm{in}}$ vs. $D_{\textrm{out}}$
Table 1: Summary of notations
@@
Figure 1 summarizes the end-to-end PANORAMIA privacy audit while Table 1 summarizes the notations used in the rest of the paper. The privacy audit starts with a target model $f$, and a subset of its training data $D_{f}$ from distribution ${\mathcal{D}}$. For instance, ${\mathcal{D}}$ and $D_{f}$ could be the distribution and dataset of one participant in an FL training procedure which returns a final model $f$. The audit then proceeds in two phases.
Phase 1: In the first phase, PANORAMIA uses a subset of the known training data $D_{G}\subset D_{f}$ to train a generative model ${\mathcal{G}}$. The goal of the generative model ${\mathcal{G}}$ is to match the training data distribution ${\mathcal{D}}$ as closely as possible, which we formalize in Definition 34). At a high level, we require the generative model to put a high likelihood on data that is also likely under ${\mathcal{D}}$. This means that the generator needs to generate a large number of data points that are good (likely under ${\mathcal{D}}$), though it is allowed to occasionally generate bad samples (unlikely under ${\mathcal{D}}$). In §4, we demonstrate that this one-sided requirement is sufficient due to our focus on finding members (i.e., true positives) as opposed to also finding non-members (i.e., true negatives) in our audit game. PANORAMIA is agnostic to the generative modeling approach, as the only primitive we need is that of data synthesis, with potential approaches including GANs, VAEs or Diffusion Models. We can also fine-tune existing generators and filter generated samples as a post-processing phase. We leverage several such approaches in practice (§5, Appendix B). Using the generative model ${\mathcal{G}}$, we can synthesize non-member data, which corresponds to data that was not used in the training of target model $f$. Hence, we now have access to an independent dataset of member data $D_{\textrm{in}}=D_{f}\backslash D_{G}$, and a synthesized dataset of non-member data $D_{\textrm{out}}$ from ${\mathcal{G}}$, that we choose to be of the same size $m$ as $D_{\textrm{in}}$.
@@
Phase 2: In the second phase, we leverage $D_{\textrm{in}}$ and $D_{\textrm{out}}$ to audit the privacy leakage of $f$ using a MIA. To this end, we split $D_{\textrm{in}},D_{\textrm{out}}$ into training and testing sets, respectively called $D^{\textrm{tr}}_{\textrm{in}},D^{\textrm{tr}}_{\textrm{out}}$ and $D^{\textrm{te}}_{\textrm{in}},D^{\textrm{te}}_{\textrm{out}}$. We use the training set to train a MIA (called PANORAMIA in Figure 1), a binary classifier that predicts whether a given datapoint is a member of $D_{f}$, the training set of the target model $f$. This MIA classifier makes its prediction based on both a training example $x$, as well as information from applying the target model $f$ to the input, such as the loss of the target model when applied to this example $\textrm{loss}(f(x))$, or the likelihood of the data $f(x)$ for language models (see §5, Appendix B for details). As discussed in §2, MIA performance is directly related to a lower-bound on possible $\epsilon$-DP values for the target model $f$. We use the test set to measure the MIA performance, using the precision at different recall values.
@@
Previous results linking the performance of a MIA on several data-points to $\epsilon$-DP bounds rely on independence between members and non-members. This intuitively means that there is no information about membership in $x$. When the auditor controls the training process this independence is enforced by construction, by adding data points to the training set based on an independent coin flip. In PANORAMIA, we do not have independence between membership and $x$, as all non-members come from the generator ${\mathcal{G}}\neq{\mathcal{D}}$. Intuitively, this implies that there are two ways to guess membership and have high MIA precision: either using $f$ to detect membership (i.e., symptomatic of privacy leakage) or detecting generated data (i.e., not a symptom of privacy leakage). Since we are only interested in measuring privacy leakage, we cannot compare our MIA results to that of a random guess but have to account for the detection of generated data based on $x$ only. To do this, we compare the results of the MIA to that of a baseline classifier $b$ that guesses membership based exclusively on $x$, without access to $f$ or its predictions on $x$. The stronger this baseline, the better the removal of the effect of synthesized data detection. §5.1 details how to construct a strong baseline while Algorithm 1 summarizes the entire procedure. In the next section, we demonstrate how to relate the difference between the baseline $b$ and the MIA performance to the privacy loss $\epsilon$.

### v2

@@ Abstract
We present PANORAMIA, a privacy leakage measurement framework for machine learning models that relies on membership inference attacks using generated data as non-members. By relying on generated non-member data, PANORAMIA eliminates the common dependency of privacy measurement tools on in-distribution non-member data. As a result, PANORAMIA does not modify the model, training data, or training process, and only requires access to a subset of the training data. We evaluate PANORAMIA on ML models for image and tabular data classification, as well as on large-scale language models.
@@ 1 Introduction
Training Machine Learning (ML) models with Differential Privacy (DP)  Dwork et al. (2006), such as with DP-SGD Abadi et al. (2016), upper-bounds the worst-case privacy loss incurred by the training data. In contrast, privacy auditing aims to empirically lower-bound the privacy loss of a target ML model or algorithm. In practice, privacy audits usually rely on the link between DP and the performance of membership inference attacks (MIA)  Wasserman & Zhou (2010); Kairouz et al. (2015); Dong et al. (2019). At a high level, DP implies an upper-bound on the performance of MIAs, thus creating a high-performance MIA implies a lower-bound on the privacy loss. Auditing schemes have proven valuable in many settings, such as to audit DP implementations Nasr et al. (2023), or to study the tightness of DP algorithms Nasr et al. (2021); Lu et al. (2023); Steinke et al. (2023). Typical privacy audits rely on retraining the model several times, each time guessing the membership of one sample Jagielski et al. (2020); Carlini et al. (2022a); Zanella-Béguelin et al. (2022), which is computationally prohibitive, requires access to the target model (entire) training data as well as control over the training pipeline.
To circumvent these concerns, Steinke et al. (2023) proposed an auditing recipe (called O(1)) requiring only one training run (which could be the same as the actual training) by randomly including/excluding several samples (called auditing examples) into the training dataset of the target model. Later, the membership of the auditing examples are guessed for privacy audit. However, O(1) faces a few challenges in certain setups. First, canaries, which are datapoints specially crafted to be easy to detect when added to the training set Nasr et al. (2023); Lu et al. (2023); Steinke et al. (2023), cannot be employed as auditing examples when measuring the privacy leakage for data that a contributor actually puts into the model, and not a worst case data point.
This matches a setting in which individual data contributors (e.g., a hospital in a cross-site Federated Learning (FL) setting or a user of a service that trains ML models on users’ data) measure the leakage of their own (i.e., known) partial training data in the final trained model. Second, O(1) also relies on the withdrawal of real data from the model to construct non-member in-distribution data. This is problematic in situations in which ML model owners need to conduct post-hoc audits, in which case it is too late for removal  Negoescu et al. (2023). Moreover, in-distribution audits require much more data, thus withholding many data points (typically more than the test set size) and reducing model utility. This brings us to the question: Given an instance of a machine learning model as a target, can we perform post-hoc estimation of the privacy loss with regards to a known member subset of the target model training dataset?
@@
Our contributions. We propose PANORAMIA, a new scheme for Privacy Auditing with NO Retraining by using Artificial data for Membership Inference Attacks. More precisely, we consider an auditor with access to a subset of the training data and introduce a new alternative for accessing non-members: using synthetic datapoints from a generative model trained on the member data, unlocking the limit on non-member data. PANORAMIA uses this generated data, together with known members, to train and evaluate a MIA attack on the target model to audit (§3). We also adapt the theory of privacy audits, and show how PANORAMIA can estimate the privacy loss (though not a lower-bound) of the target model with regards to the known member subset (§4). An important benefit of PANORAMIA is to perform privacy loss measurements with (1) no retraining the target ML model (i.e., we audit the end-model, not the training algorithm), (2) no alteration of the model, dataset, or training procedure, and (3) only partial knowledge of the training set. We evaluate PANORAMIA on CIFAR10 models and observe that overfitted models, larger models, and models with larger DP parameters have higher measured privacy leakage. We also demonstrate the applicability of our approach on the GPT-2 based model (i.e., WikiText dataset) and CelebA models.
2 Background and Related Work
@@
DP is the established privacy definition in the context of ML models, as well as for data analysis in general. We focus on the pure DP definition to quantify privacy loss with well-understood semantics. In a nutshell, DP is a property of a randomized mechanism (or computation) from datasets to an output space ${\mathcal{O}}$, noted $M:{\mathcal{D}}\rightarrow{\mathcal{O}}$. It is defined over neighboring datasets $D,D^{\prime}$, differing by one element $x\in{\mathcal{X}}$ (we use the add/remove neighboring definition), which is $D^{\prime}=D\cup\{x\}$. Formally:
@@
A mechanism $M:{\mathcal{D}}\rightarrow{\mathcal{O}}$ is $\epsilon$-DP if for any two neighbouring datasets $D$, $D^{\prime}\in{\mathcal{D}}$, and for any measurable output subset $O\subseteq{\mathcal{O}}$ it holds that:
$P[M(D)\in O]\leq e^{\varepsilon}P[M(D^{\prime})\in O].$
@@
Since the neighbouring definition is symmetric, so is the DP definition, and we also have that $P[M(D^{\prime})\in O]\leq e^{\varepsilon}P[M(D)\in O]$. Intuitively, $\epsilon$ upper-bounds the worst-case contribution of any individual example to the distribution over outputs of the computation (i.e., the ML model learned). More formally, $\epsilon$ is an upper-bound on the privacy loss incurred by observing an output $o$, defined as $\big{|}\ln\left(\frac{{\mathbb{P}}[M(D)=o]}{{\mathbb{P}}[M(D^{\prime})=o]}\right)\big{|}$, which quantifies how much an adversary can learn to distinguish $D$ and $D^{\prime}$ based on observing output $o$ from $M$. A smaller $\epsilon$ hence means higher privacy.
@@
DP, MIA and privacy audits. To audit a DP training algorithm $M$ that outputs a model $f$, one can perform a MIA on datapoint $x$, trying to distinguish between a neighboring training sets $D$ and $D^{\prime}=D\cup\{x\}$. The MIA can be formalized as a hypothesis test to distinguish between ${\mathcal{H}}_{0}=D$ and ${\mathcal{H}}_{1}=D^{\prime}$ using the output of the computation $f$. Wasserman & Zhou (2010); Kairouz et al. (2015); Dong et al. (2019) show that any such test at significance level $\alpha$ (False Positive Rate or FPR) has power (True Positive Rate or TPR) bounded by $e^{\epsilon}\alpha$. In practice, one repeats the process of training model $f$ with and without $x$ in the training set, and uses a MIA to guess whether $x$ was included. If the MIA has $\textrm{TPR}>e^{\epsilon}\textrm{FPR}$, the training procedure that outputs $f$ is not $\epsilon$-DP. This is the building block of most privacy audits Jagielski et al. (2020); Nasr et al. (2021); Zanella-Béguelin et al. (2022); Lu et al. (2023); Nasr et al. (2023). Appendix F provides further context, and discusses other related work that is not directly relevant to our contribution.
@@
Averaging over data instead of models with O(1). The above result bounds the success rate of MIAs when performed over several retrained models, on two alternative datasets $D$ and $D^{\prime}$. Steinke et al. (2023) show that it is possible to average over data when several data points independently differ between $D$ and $D^{\prime}$. Let $x_{1,m}$ be the $m$ data points independently included in the training set, and $s_{1,m}\in\{0,1\}^{m}$ be the vector encoding inclusion. $T_{0,m}\in{\mathbb{R}}^{m}$ represents any vector of guesses, with positive values for inclusion in the training set (member), negative values for non-member, and zero for abstaining. Then, if the training procedure is $\epsilon$-DP, Proposition 5.1 in Steinke et al. (2023) bounds the performance of guesses from $T$ with:
@@
${\mathbb{P}}\Big{[}\sum_{i=1}^{m}$ $\max\{0,T_{i}\cdot S_{i}\}\geq v\ |\ T=t\Big{]}\leq\underset{{S^{\prime}\sim\textrm{Bernoulli}(\frac{e^{\epsilon}}{1+e^{\epsilon}})^{m}}}{{\mathbb{P}}}\Big{[}\sum_{i=1}^{m}|t_{i}|\cdot S^{\prime}_{i}\geq v\Big{]}.$
@@
In other words, an audit (MIA) $T$ that can guess membership better than a Bernoulli random variable with probability $\frac{e^{\epsilon}}{1+e^{\epsilon}}$ refutes an $\epsilon$-DP claim. In this work we build on this result, extending the algorithm (§3) and theoretical analysis (§4) to enable the use of generated data for non-members, which breaks the independence between data points.
@@ 3 PANORAMIA
Figure 1 summarizes the end-to-end PANORAMIA privacy measurement. The measurement starts with a target model $f$, and a subset of its training data $D_{f}$ from distribution ${\mathcal{D}}$. For instance, ${\mathcal{D}}$ and $D_{f}$ could be the distribution and dataset of one participant in an FL training procedure that outputs a final model $f$. The privacy measurement then proceeds in two phases.
Figure 1: PANORAMIA’s two-phase privacy audit. Phase 1 trains generative model ${\mathcal{G}}$ on member data. Phase 2 trains a MIA on a subset of member data and generated non-member data, using the loss of $f$ on these data points. The performance of the MIA is compared to a baseline classifier that does not have access to $f$. Notations are summarized in Table 4 in Appendix A.
Phase 1: In the first phase, PANORAMIA uses a subset of the known training data $D_{G}\subset D_{f}$ to train a generative model ${\mathcal{G}}$. The goal of the generative model ${\mathcal{G}}$ is to match the training data distribution ${\mathcal{D}}$ as closely as possible, which is formalized in Definition 34). Using the generative model ${\mathcal{G}}$, we can synthesize non-member data, which corresponds to data that was not used in the training of target model $f$. Hence, we now have access to an independent dataset of member data $D_{\textrm{in}}=D_{f}\backslash D_{G}$, and a synthesized dataset of non-member data $D_{\textrm{out}}\sim{\mathcal{G}}$, of size $m=|D_{\textrm{in}}|$.
@@
Phase 2: In the second phase, we leverage $D_{\textrm{in}}$ and $D_{\textrm{out}}$ to audit the privacy leakage of $f$ using a MIA. To this end, we split $D_{\textrm{in}},D_{\textrm{out}}$ into training and testing sets, respectively called $D^{\textrm{tr}}_{\textrm{in}},D^{\textrm{tr}}_{\textrm{out}}$ and $D^{\textrm{te}}_{\textrm{in}},D^{\textrm{te}}_{\textrm{out}}$. We use the training set to train a MIA (called PANORAMIA in Figure 1), a binary classifier that predicts whether a given datapoint is a member of $D_{f}$, the training set of the target model $f$. This MIA classifier makes its prediction based on both a training example $x$, as well as information from applying the target model $f$ to the input, such as the loss of the target model when applied to this example $\textrm{loss}(f(x))$ (see §5, Appendix C for details). We use the test set to measure the MIA performance, using the precision at different recall values.
@@
Previous results linking the performance of a MIA on several data-points to $\epsilon$-DP bounds rely on independence between members and non-members. This intuitively means that there is no information about membership in $x$ itself. When the auditor controls the training process this independence is enforced by construction, by adding data points to the training set based on an independent coin flip. In PANORAMIA, we do not have independence between membership and $x$, as all non-members come from the generator ${\mathcal{G}}\neq{\mathcal{D}}$. As a result, there are two ways to guess membership and have high MIA precision: either by using $f$ to detect membership (i.e., symptomatic of privacy leakage) or by detecting generated data (i.e., not a symptom of privacy leakage). To measure the privacy leakage, we compare the results of the MIA to that of a baseline classifier $b$ that guesses membership based exclusively on $x$, without access to $f$. The stronger this baseline, the better the removal of the effect of synthesized data detection. Algorithm 1 summarizes the entire procedure. In the next section, we demonstrate how to relate the difference between the baseline $b$ and the MIA performance to the privacy loss $\epsilon$.